A roller coaster car changes velocity from 2.5 m/s to 13.3 m/s as it rolls down a hill for 3 seconds. What is the car’s acceleration?
1
Expert's answer
2019-06-06T10:05:23-0400

Assuming the car's acceleration to be constant, the change of the velocity can be calculated as follows:


"v(t) = v_0 + a t,"

where a is the acceleration of a car.

Hence,


"a = \frac{v-v_0}{t}"

Substituting the numerical values, we obtain:


"a = \frac{13.3 - 2.5}{3} =3.6 \, m/s^2"

Answer: 3.6 m/s2